# WikiGap: Promoting Epistemic Equity by Surfacing Knowledge Gaps Between English Wikipedia and other Language Editions

### v2

@@ Abstract.
With more than 11 times as many pageviews as the next, English Wikipedia dominates global knowledge access relative to other language editions. Readers are prone to assuming English Wikipedia as a superset of all language editions, leading many to prefer it even when their primary language is not English. Other language editions, however, comprise complementary facts rooted in their respective cultures and media environments, which are marginalized in English Wikipedia. While Wikipedia’s user interface enables switching between language editions through its Interlanguage Link (ILL) system, it does not reveal to readers that other language editions contain valuable, complementary information. We present WikiGap, a system that surfaces complementary facts sourced from other Wikipedias within the English Wikipedia interface. Specifically, by combining a recent multilingual information-gap discovery method (Samir et al., 2024) with a user-centered design, WikiGap enables access to complementary information from French, Russian, and Chinese Wikipedia. In a mixed-methods study (n=21), WikiGap significantly improved fact-finding accuracy, reduced task time, and received a 32-point higher usability score relative to Wikipedia’s current ILL-based navigation system. Participants reported increased awareness of the availability of complementary information in non-English editions and reconsidered the completeness of English Wikipedia. WikiGap thus paves the way for improved epistemic equity across language editions.
@@ 1. Introduction
In practice, however, the overwhelming dominance of English poses a significant challenge to egalitarian knowledge curation. It is not uncommon for such editors to contribute culture- or region-specific knowledge to English Wikipedia to encounter hegemonic resistance. While Wikipedia is in principle open to contributions from any person with literacy skills and an internet connection, it was reported in (2017) that an overwhelming majority of edits are contributed by a small portion of the editor base (Oberhaus, 2017). As of (2024), for English Wikipedia, these editors were overwhelmingly male and white (Wikimedia Foundation, 2024). This has resulted in contributions being rejected, even when the topic is geographically positioned in a non-Anglophone country; see, for example, a case study on a lengthy editor dispute resulting in the discarding of Indian editors’ cultural knowledge of the Ganga river in India (Kumar, 2017).
Yet, at the same time, English Wikipedia is by far the most widely read and authoritative source among the $300+$ language editions. In a 2011 report, $93\%$ of all editors surveyed attested to reading English Wikipedia, even though it was the primary language for only $52\%$ of them (Kumar, 2021; Wikimedia Foundation, 2011). We are thus presented with a remarkable asymmetry – English Wikipedia marginalizes potential contributions from non-Anglophone sources, yet inadvertently monopolizes attention from information-seeking readers away from language editions that contain these marginalized perspectives (e.g., Avieson and Mazumdar, 2025). This epistemic assumption of the completeness of English Wikipedia, which gives rise to this asymmetrically large readership, was previously aptly named the English-as-superset hypothesis (Hecht, 2013).
@@ 2. Background and Related Work
Wikipedia is a peer production system (Benkler et al., 2015), where a large number of collaborators come together to produce a shared outcome. There are important norms in this collaborative practice that are referenced heavily in deliberating content contributions, although the explication and invocation of these norms can vary widely across language editions (Hwang and Shaw, 2022; Bipat et al., 2021). It has also introduced mechanisms like page protection policies, which increase friction in contributing edits to popular, controversial pages (Ajmani et al., 2023b). But the overarching ethos is that there should be no hard-set rules, and Wikipedia should remain largely self-organizing. Large-scale collaborative peer production systems like Wikipedia are intended to democratize access to information, making it “easily available to people who would have otherwise had to seek it out in piecemeal and often haphazard fashion, if they could access it at all” (Justice and Golumbia, 2024, pp. 273). It is not simply about access either; Wikipedia is largely associated with an air of democratizing the production of knowledge, a telos of rebelling against centralized authority and bureaucracy. This more laissez-faire form of collaborative content production is underpinned by the idea that a distributed form of knowledge curation could result in egalitarian outcomes in the access to and production of knowledge.
@@
But this vision has been argued to be highly idealistic, as the capacity to create archival encyclopedia content has long been recognized as sociopolitically contingent (Foucault, 2013; Said, 1977). For example, the composition of English Wikipedia editor base (white, male) has predictably led to the marginalization of women on Wikipedia (Hargittai and Shaw, 2015), as well as scholars from racial minority backgrounds (Adams et al., 2019). When it comes to multilingualism in Wikipedia, Wikipedia’s pillar for verifiable information, a component of the Neutral Point of View policy, has introduced barriers for creating content in other language editions. English is widely recognized as the preeminent vehicle for conveying scholarship, making it challenging to find primary sources for other language editions unless editors are fluent in English (Sen et al., 2015). While there are multilingual editors who contribute to multiple language editions, who serve as a bridge connecting different language editions (Hickman et al., 2021), they are a smaller proportion of editors on Wikipedia (Hale, 2014).
@@
In this section, we highlight the perceived hegemonic status of the English language edition, while maintaining that it has information gaps that illustrate that it should not serve as the authoritative language edition (Sec. 2.1). We then discuss previous solutions, including a unified language edition, Wikipedia’s Interlanguage Link navigation panel, and other interfaces proposed for interacting with content in different language editions (Sec. 2.2). We argue that these prior works have yet to achieve an important goal, increasing readers’ interaction with content from other language editions. In achieving this goal, we could challenge the perceived completeness of the English language edition, what Hecht (2013) has previously called the English-as-superset hypothesis, an important step towards greater democratization of knowledge production. In Sec. 3, we introduce our system, WikiGap, that augments the experience of reading English Wikipedia documents by showcasing information gaps relative to other language editions.
@@ 2.1. English Wikipedia is Considered the Default, yet Information Gaps with Other Editions Persist
While some facts are culturally invariant, such as physical laws, some of Wikipedia’s most-read articles pertain to socio-geographically positioned histories of peoples, places, recipes, and so on. Generally speaking, all of Wikipedia’s language editions are necessarily socio-geographically positioned (Johnson and Lescak, 2022) and exhibit a “self-focus bias(Hecht and Gergle, 2009), the tendency of editors to contribute more to articles relevant to their region, including articles on people (Samir et al., 2024; Callahan and Herring, 2011), places (Oeberst and Ridderbecks, 2024), and cuisines (Laufer et al., 2015).
@@
For those culturally-positioned topics, the presupposition of a universal language-agnostic knowledge base has contributed to the marginalization of other language editions. English Wikipedia is treated as the default, standard, or “invisible center(Benjamin, 2023) through which other language editions are compared against. For example, Hickman et al. (2021) documents that bilingual Indian and Pakistani editors refer to English Wikipedia to determine the contributions to make to their local language editions, perceiving English Wikipedia to be the objective standard. [^2] This isn’t exclusive to South Asian editors; as we show in Section LABEL:section:evaluation, readers and editors alike remain largely unaware of cross-linguistic heterogeneity in Wikipedia. In reality, neither the English version nor the local language edition for a topic is a complete article; a considerable proportion of unique facts can still be identified in non-local language editions (Samir et al., 2024). There is thus an incentive to look at multiple language editions, even when a topic has a strong regional association with one language edition.
@@ 2.2. Prior Work Targeting Cross-Lingual Knowledge Dissemination
Currently, Wikipedia supports accessing different language editions’ coverage of topics by leveraging Interlanguage Links. These links can be accessed through a dropdown box while reading any given Wikipedia article. While readers can ostensibly leverage the hyperlinks in this dropdown box to gain access to complementary information in different language editions, this panel fails to indicate to readers that language editions often contain large proportions of complementary content. Rather, the hyperlinks suggest a change in linguistic encoding, akin to changing languages in a system like Google Translate, rather than wholesale changes in composition and curation. Indeed, in our user studies (Section 5), we find that readers typically believe that language editions share an underlying knowledge base.
@@ 3.1.1. Design Requirements Informed by User Needs and Theory
Instead, users preferred lightweight, non-intrusive cues – similar to annotations in Grammarly or Google Docs – that subtly signal the presence of additional information (F2). This aligns with Anchored Annotation Theory (Marshall, 1997), which finds that tying annotations to specific text locations improves comprehension and recall. Our design anchors cross-lingual facts directly to the English article using hover-triggered underlines and a collapsible sidebar, allowing users to access additional facts only when interested. This approach also follows the Spatial Contiguity Principle (Johnson and Mayer, 2012), which emphasizes placing related information close together. Together, these design choices preserve reading flow while enabling low-friction, in-place access to multilingual content.
@@
To meet this need, we prioritize presenting each fact with an explicit language label, consistently color-coded across the interface. Additionally, readers should be able to open the original article and view the corresponding sentence in context. These strategies support transparency and accountability, giving readers confidence in the surfaced content and encouraging deeper multilingual exploration.
@@ 3.1.2. Core Design Elements
D1: Cross-Lingual Gap Anchors.
To address R1 (non-disruptive presentation) and R2 (traceability), we implemented a subtle underlining system that indicates the presence of additional information from other language editions. These highlights use an underline pattern in color to indicate availability without disrupting reading flow. When hovered or clicked, they become slightly more prominent to provide interaction feedback and invite further exploration. Each language is associated with a distinct color – red for Chinese, blue for French, and green for Russian. This color coding is consistently applied across both in-text highlights and corresponding elements in the sidebar, reinforcing source attribution as emphasized in R2. As we elaborate in Sec. 3.3.2, only significant knowledge gaps identified by InfoGap are highlighted, rather than minor variations or differences in wording. This selective highlighting prevents information overload and ensures that users’ attention is drawn to substantive cross-lingual differences.
D2: Cross-Lingual Fact Sidebar
@@
D3: Source Language Toggle.
To support R3 (customizable exploration), we introduced a language filter function located at the top of the sidebar. These filters allow users to specify which language(s) they want to view content from, enabling personalization based on their linguistic background and interests. The filters use both language names and visual indicators (e.g., country flag and colors) to maximize clarity. This functionality directly supports both passive and active interaction modes (R3).
D4: Cross-Lingual Fact Cards.
@@
D5: Multilingual Fact Search.
@@ 3.2. The InfoGap Pipeline
Figure 3. Knowledge differences in the Wikipedia coverage of Oolong identified by InfoGap. Connecting lines mark overlapping facts; green boxes highlight facts unique to each language edition.
@@ 3.3. Extending and Integrating InfoGap for WikiGap
The original InfoGap pipeline was developed for detecting cross-lingual gaps in biographical articles between English and either Russian or French. To generalize the method to a new domain and an additional language, we incorporated support for Chinese into the fact decomposition and verification modules. This involved custom preprocessing for tokenizing Chinese text, crafting prompts in Chinese, and performing human evaluation on Chinese articles to ensure that InfoGap is as reliable for Chinese as it is for the original languages studied by Samir et al. (2024).
@@ 3.3.2. Language attribution and fact selection
Because WikiGap displays cross-lingual content within the English Wikipedia page to readers who we do not assume read other languages, we translate the gap facts from their original language ($L_{t}$) into English for presentation in the fact cards. To support source traceability and deeper exploration, we create a direct link-to-highlight for each fact. An encoded version of the original sentence in $L_{t}$ is appended to the target article’s URL, sending readers to the exact sentence in context when they click on the “View on [$L_{t}$] Wikipedia” in the card (D4).
@@ 4. Evaluation Methods
Table 1. Description of Study Metrics and Corresponding Data Collection and Analysis
@@
Quiz Accuracy Percentage of correct answers per condition Paired t-test comparing WikiGap vs. control conditions
@@
Usability System Usability Scale (SUS) score per condition Paired t-test comparing SUS scores between conditions
@@ 4.1. Participants
Twenty-one people participated in the study. Of these, eight were student volunteers from the HCI course where this project originated, and the remaining thirteen participants were recruited through Upwork, and were compensated at a rate of 20 USD for the user study. Participants ranged in age from 18 to 44 and represented diverse cultural and ethnic backgrounds (see Figure 6). All participants reported that they typically use the English version of Wikipedia. Twelve had previously contributed to Wikipedia to varying degrees, with two considering themselves frequent editors. Participants were required to meet the following inclusion criteria: (i) regular use of English Wikipedia, (ii) comfort with using a Chrome browser, and (iii) willingness to contribute user experience feedback.
The study followed a within-subject design, in which each participant was assigned two out of five topics (Sec. 4.3). Each participant read one article using the WikiGap extension and another article using the default Wikipedia interface as a control, in a randomized order. The assignment of treatment and control conditions was randomized for each participant.
Figure 6. Demographics of Participants.

### v3

@@ Abstract.
With more than 11 times as many pageviews as the next largest edition, English Wikipedia dominates global knowledge access relative to other language editions. Readers are prone to assuming English Wikipedia as a superset of all language editions, leading many to prefer it even when their primary language is not English. Other language editions, however, comprise complementary facts rooted in their respective cultures and media environments, which are marginalized in English Wikipedia. While Wikipedia’s user interface enables switching between language editions through its Interlanguage Link (ILL) system, it does not reveal to readers that other language editions contain valuable, complementary information. We present WikiGap, a system that surfaces complementary facts sourced from other Wikipedias within the English Wikipedia interface. Specifically, by combining a recent multilingual information-gap discovery method (Samir et al., 2024) with a user-centered design, WikiGap enables access to complementary information from French, Russian, and Chinese Wikipedia. In a mixed-methods study (n=21), WikiGap significantly improved fact-finding accuracy, reduced task time, and received a 32-point higher usability score relative to Wikipedia’s current ILL-based navigation system. Participants reported increased awareness of the availability of complementary information in non-English editions and reconsidered the completeness of English Wikipedia. WikiGap thus paves the way for improved epistemic equity across language editions.
@@ 1. Introduction
In practice, however, the overwhelming dominance of English poses a significant challenge to egalitarian knowledge curation. It is not uncommon for such editors to contribute culture- or region-specific knowledge to English Wikipedia to encounter hegemonic resistance. While Wikipedia is in principle open to contributions from any person with literacy skills and an internet connection, it was reported in (2017) that an overwhelming majority of edits are contributed by a small portion of the editor base (Oberhaus, 2017). As of (2024b), for English Wikipedia, these editors were overwhelmingly male and white (Wikimedia Foundation, 2024b). This has resulted in contributions being rejected, even when the topic is geographically positioned in a non-Anglophone country; see, for example, a case study on a lengthy editor dispute resulting in the discarding of Indian editors’ cultural knowledge of the Ganga river in India (Kumar, 2017).
Yet, at the same time, English Wikipedia is by far the most widely read and authoritative source among the $300+$ language editions. In a 2011 report, $93\%$ of all editors surveyed attested to reading English Wikipedia, even though it was a primary language for only $52\%$ of them (Kumar, 2021; Wikimedia Foundation, 2011). We are thus presented with a remarkable asymmetry – English Wikipedia marginalizes potential contributions from non-Anglophone sources, yet inadvertently monopolizes attention from information-seeking readers away from language editions that contain these marginalized perspectives (e.g., Avieson and Mazumdar, 2025). This epistemic assumption of the completeness of English Wikipedia, which gives rise to this asymmetrically large readership, was previously aptly named the English-as-superset hypothesis (Hecht, 2013).
@@ 2. Background and Related Work
Wikipedia is a peer production system (Benkler et al., 2015), where a large number of collaborators come together to produce a shared outcome. There are important norms in this collaborative practice that are referenced heavily in deliberating content contributions, although the explication and invocation of these norms can vary widely across language editions (Hwang and Shaw, 2022; Bipat et al., 2021). Community-developed tools such as vandalism patrol and bot-assisted help sustain these collaborative norms (Geiger and Ribes, 2010). Wikipedia has also introduced formal mechanisms like page protection policies, which increase friction in contributing edits to popular or controversial pages (Ajmani et al., 2023b). But the overarching ethos is that there should be no hard-set rules, and Wikipedia should remain largely self-organizing. These tensions between coordination, conflict, and self-organization mirror challenges seen across peer-production communities beyond Wikipedia, in thousands of other online wiki communities, for instance Wookiepedia and the Psychology Wiki (Kittur and Kraut, 2010). Large-scale collaborative peer production systems like Wikipedia are intended to democratize access to information, making it “easily available to people who would have otherwise had to seek it out in piecemeal and often haphazard fashion, if they could access it at all” (Justice and Golumbia, 2024, pp. 273). It is not simply about access either; Wikipedia is largely associated with an air of democratizing the production of knowledge, a telos of rebelling against centralized authority and bureaucracy. This more laissez-faire form of collaborative content production is underpinned by the idea that a distributed form of knowledge curation could result in egalitarian outcomes in the access to and production of knowledge.
But this vision has been argued to be highly idealistic, as the capacity to create archival encyclopedia content has long been recognized as sociopolitically contingent (Foucault, 2013; Said, 1977). For example, the composition of the English Wikipedia editor base (white, male) has predictably led to the marginalization of women on Wikipedia (Hargittai and Shaw, 2015; Wagner et al., 2016), as well as scholars from racial minority backgrounds (Adams et al., 2019). The demographic skews in the editor base are not merely incidental, but they can systematically shape whose knowledge is valued in the encyclopedia, privileging some forms of knowledge while marginalizing others. This can lead to reproducing epistemic injustice (Ajmani et al., 2024).
@@
When it comes to multilingualism in Wikipedia, Wikipedia’s pillar for verifiable information, a component of the Neutral Point of View policy, has introduced barriers for creating content in other language editions. English is widely recognized as the preeminent vehicle for conveying scholarship, making it challenging to find primary sources for other language editions unless editors are fluent in English (Sen et al., 2015). While there are multilingual editors who contribute to multiple language editions, who serve as a bridge connecting different language editions (Hickman et al., 2021), they are a smaller proportion of editors on Wikipedia (Hale, 2014).
@@
In this section, we highlight the perceived hegemonic status of the English language edition, while maintaining that it has information gaps that illustrate that it should not serve as the authoritative language edition (Sec. 2.1). We then discuss previous solutions, including a unified language edition, Wikipedia’s Interlanguage Link navigation panel, and other interfaces proposed for interacting with content in different language editions (Sec. 2.2). We argue that these prior works have yet to achieve an important goal of increasing readers’ interaction with content from other language editions. In achieving this goal, we challenge the perceived completeness of the English language edition—what Hecht (2013) has previously called the English-as-superset hypothesis—an important step towards greater democratization of knowledge production. In Sec. 3, we introduce our system, WikiGap, that augments the experience of reading English Wikipedia documents by showcasing information gaps relative to other language editions.
@@ 2.1. English Wikipedia is Considered the Default, yet Information Gaps with Other Editions Persist
While some facts are culturally invariant, such as physical laws, some of Wikipedia’s most-read articles pertain to socio-geographically positioned histories of peoples, places, recipes, and so on (Wikimedia Foundation, 2024a). Generally speaking, all of Wikipedia’s language editions are necessarily socio-geographically positioned (Johnson and Lescak, 2022) and exhibit a “self-focus bias(Hecht and Gergle, 2009), the tendency of editors to contribute more to articles relevant to their region, including articles on people (Samir et al., 2024; Callahan and Herring, 2011), places (Oeberst and Ridderbecks, 2024), and cuisines (Laufer et al., 2015).
@@
For those culturally-positioned topics, the presupposition of a universal language-agnostic knowledge base has contributed to the marginalization of other language editions. English Wikipedia is treated as the default, standard, or “invisible center(Benjamin, 2023) through which other language editions are compared against. For example, Hickman et al. (2021) documents that bilingual Indian and Pakistani editors refer to English Wikipedia to determine the contributions to make to their local language editions, perceiving English Wikipedia to be the objective standard. [^2] This isn’t exclusive to South Asian editors; as we show in Section 5, readers and editors alike remain largely unaware of cross-linguistic heterogeneity in Wikipedia. In reality, neither the English version nor the local language edition for a topic is a complete article; a considerable proportion of unique facts can still be identified in non-local language editions (Samir et al., 2024). There is thus an incentive to look at multiple language editions, even when a topic has a strong regional association with one language edition.
@@ 2.2. Prior Work Targeting Cross-Lingual Knowledge Dissemination
Currently, Wikipedia supports accessing different language editions of a topic through Interlanguage Links, available via a dropdown box on each article. While these links ostensibly provide readers with access to complementary information across editions, the interface gives little indication that articles may differ substantially in content. Instead, the hyperlinks imply only a change in linguistic encoding, similar to using Google Translate, and do not signal the significant differences in composition and curation across editions. Consistent with this design, our user studies (Section 5) show that many readers assume language editions share a single underlying knowledge base.
@@ 3.1.1. Design Requirements Informed by User Needs and Theory
Instead, users preferred lightweight, non-intrusive cues – similar to annotations in Grammarly or Google Docs – that subtly signal the presence of additional information (F2). This aligns with Anchored Annotation Theory (Marshall, 1997), which finds that tying annotations to specific text locations improves comprehension and recall. Our design anchors cross-lingual facts directly to the English article using click-triggered underlines and a collapsible panel, allowing users to access additional facts only when interested. This approach follows the Spatial Contiguity Principle (Johnson and Mayer, 2012), which emphasizes placing related information close together. Collectively, these design choices preserve reading flow while enabling low-friction, in-place access to multilingual content.
@@
To meet this need, we present each fact with an explicit, color-coded language label and allow readers to trace each fact back to its source by opening the corresponding language edition and viewing the sentence in context. These strategies support transparency and accountability, giving readers confidence in the surfaced content and encouraging deeper multilingual exploration.
@@ 3.1.2. Core Design Elements
D1: In-text Gap Markers.
To address R1 (non-disruptive presentation) and R2 (traceability), we implemented a subtle underlining system that signals the presence of additional information from other language editions. Each marker is country-color coded and appears as a patterned underline within the English article. When clicked, the side panel retrieves the fact relevant to that sentence, where the extension identifies complementary or missing information from another language edition. The underline then becomes more prominent, providing interaction feedback and inviting further exploration. Each language is associated with a distinct color – red for Chinese, blue for French, and green for Russian. This color coding is applied across both in-text highlights and corresponding elements in the side panel, reinforcing source attribution as emphasized in R2. As we elaborate in Sec. 3.3.2, only significant knowledge gaps identified by InfoGap are highlighted, rather than minor variations or differences in wording. This selective highlighting prevents information overload and ensures that users’ attention is drawn to substantive cross-lingual differences.
D2: Cross-Lingual Fact Panel
@@
D3: Fact Provenance Card
@@
D4: Source Language Filter
To support R3 (customizable exploration), we introduced a language filter function located at the top of the sidebar. These filters allow users to specify which language(s) they want to view content from, enabling personalization based on their linguistic background and interests. The filters use both language names and colors to maximize clarity. This functionality directly supports both passive and active interaction modes (R3).
D5: Cross-lingual Fact Search.
@@ 3.2. The InfoGap Pipeline
Figure 3. Knowledge differences in the Wikipedia coverage of Oolong identified by InfoGap. Connecting lines mark overlapping facts; green boxes highlight facts unique to each language edition. English sentences in italics represent translations of facts from Chinese or French Wikipedia, while non-italicized English sentences are from the original English Wikipedia article.
@@ 3.3. Extending and Integrating InfoGap for WikiGap
The original InfoGap pipeline was developed for detecting cross-lingual gaps in biographical articles between English and either Russian or French. To generalize the method to a new domain and an additional language, we incorporated support for Chinese into the fact decomposition and verification modules. This involved custom preprocessing for tokenization, crafting prompts, and conducting human evaluation on Chinese texts to confirm that InfoGap performs comparably to its performance on the original languages studied by Samir et al. (2024).
@@ 3.3.2. Language attribution and fact selection
Because WikiGap displays cross-lingual content within the English Wikipedia page to readers who we do not assume read other languages, we translate the gap facts from their original language ($L_{t}$) into English for presentation in the fact cards. To support source traceability and deeper exploration, we create a direct link-to-highlight for each fact. An encoded version of the original sentence in $L_{t}$ is appended to the target article’s URL, sending readers to the exact sentence in context when they click on the “View on $L_{t}$] Wikipedia” in the card (D3).
@@ 4. Evaluation Methods
Table 1. Description of study metrics and corresponding data collection and analysis. All metrics were collected at the individual participant level and compared across WikiGap and control conditions.
@@
Quiz Accuracy Percentage of correct answers in each condition Paired t-test comparing WikiGap vs. control conditions
@@
Usability System Usability Scale (SUS) score in each condition Paired t-test comparing SUS scores between conditions
@@ 4.1. Participants
Twenty-one people participated in the study. Of these, eight were student volunteers from the HCI course where this project originated, and the remaining thirteen participants were recruited through Upwork, and were compensated at a rate of 20 USD for the user study. Participants ranged in age from 18 to 44 and represented diverse cultural and ethnic backgrounds (see Table 2). All participants reported that they typically use the English version of Wikipedia. Twelve had previously contributed to Wikipedia to varying degrees, with two considering themselves frequent editors. Participants were required to meet the following inclusion criteria: (i) regular use of English Wikipedia, (ii) comfort with using a Chrome browser, and (iii) willingness to contribute user experience feedback.
The study followed a within-subject design, in which each participant was assigned two out of five topics (Sec. 4.3). Each participant read one article using the WikiGap extension and another article using the default Wikipedia interface (with ILLs) as a control. The assignment of treatment and control conditions was randomized for each participant.
Table 2. Participant self-reported cultural background.
Cultural / Ethnic background % of participants
East Asian 24%
White or European 19%
South Asian 14%
Hispanic or Latin American 14%
Black or African 10%
Middle Eastern or North Southeast Asian 10%
African 5%
Other 5%